OpenStudy (anonymous):

A textile manufacturer has daily production costs of C(x)=10,000-110x+0.045x^2 where C is the total cost in dollars and x is the number of units produced. how many fixtures x should be produced each day to yield a minimum cost?

OpenStudy (anonymous):

Question is not complete.

OpenStudy (campbell_st):

have you studied calculus..?

OpenStudy (anonymous):

its off of a sample final for precalculus

OpenStudy (campbell_st):

ok... well here is an easy solution. the curve is a parabola... the minimum cost lies on the line of symmetry for the parabola... when given an equation \[ax^2 + bx + c = 0\] the line of symmetry is \[x = \frac{-b}{2a}\] in your question b = - 110 and a = 0.045 substitute and it will give how many fixtures need to be produced to minimise the cost to get the minimum cost, substitute you value into the original equation. hope it helps

OpenStudy (anonymous):

\[\frac{ -110 }{ 2(0.045)}\] =-1222.22, so am I supposed to plug that in for x in the original problem? because when I do I get 211666, which definitely does not seem right.

OpenStudy (anonymous):

Yep, there is something wrong with that question. Something is missing.

OpenStudy (anonymous):

Alright thanks for trying, they never made all that much since at the beginning of the semester either.